### v2

# Peng Cheng, and Jiming Chen

Abstract
Reinforcement learning (RL) is a powerful tool for sequential decision-making, but its application is often hindered by privacy concerns arising from its interaction data. This challenge is particularly acute in advanced Cyber-Physical Systems (CPS), where learning from operational and user data can expose systems to privacy inference attacks. Existing differential privacy (DP) models for RL are often inadequate: the centralized model requires a fully trusted server, creating a single point of failure risk, while the local model incurs significant performance degradation that is unsuitable for many control applications. This paper addresses this gap by leveraging the emerging shuffle model of privacy, an intermediate trust model that provides strong privacy guarantees without a centralized trust assumption. We present Shuffle Differentially Private Policy Elimination (SDP-PE), the first generic policy elimination-based algorithm for episodic RL under the shuffle model. Our method introduces a novel exponential batching schedule and a “forgetting” mechanism to balance the competing demands of privacy and learning performance. Our analysis shows that SDP-PE achieves a near-optimal regret bound, demonstrating a superior privacy-regret trade-off that significantly outperforms the local model. This work establishes the viability of the shuffle model for secure data-driven control in advanced CPS.
Index Terms:
Reinforcement Learning, Shuffle Differential Privacy, Online Learning.
I Introduction
Reinforcement learning (RL) is a sequential decision-making framework that has gained remarkable attraction for CPS, such as smart grid, networked robotics, and autonomous vehicle networks. In these systems, a learning agent continuously interacts with the environment to receive new observations and rewards, and improves the performance of its policy. However, a crucial concern for its deployment is that the interaction data usually contains sensitive information. This concern is not only limited to personal data in user-centric CPS [1], but also extends to confidential operational data, such as control parameters or system states in a distributed network [2]. Without privacy protection mechanisms in place, the learning agent can memorize sensitive information from the interaction history [3], exposing confidential operational patterns or leaving the system’s underlying entities vulnerable to various privacy attack [4].
To counter such threat, differential privacy (DP) [5] has been extensively applied as a formal privacy method in various decision-making settings, including multi-armed bandits [6], classical control systems [7, 8], and networked control systems [9, 10]. In the case of RL, a growing body of literature has addressed the privacy of different components of the learning process, such as protecting the reward function [11], the transition dynamics [12], or the learned policies [13]. Distinct from this line of work, our focus is on protecting the interaction history itself. Prior work has primarily focused on two privacy architectures: the centralized privacy model with Joint DP (JDP) constraint [14, 15] and the local privacy model with Local DP (LDP) constraint [16, 17]. To apply these frameworks, the standard approach is to consider an episodic online interaction setting. In this formulation, the agent interacts with the environment for $K$ episodes, and each episode is regarded as an interaction with one specific user. In a CPS context, this “user” could be a human customer, a smart home, or an autonomous vehicle providing data. The aim is to ensure that the user’s states and generated rewards will not be inferred by an adversary during the learning process. Specifically, JDP guarantees that all other users’ decisions will not leak much information about any specific user, and $\varepsilon$-JDP can be obtained at the expense of an additive logarithmic term of $\mathcal{O}(\log(K)/\varepsilon)$ in the regret bound [14]. However, a trusted central server is required to collect and process the raw interaction data under JDP, which may result in a single point of failure. On the other hand, $\varepsilon$-LDP removes this risk by privatizing data on each users’ local device, but at the cost of severe performance degradation, resulting in a multiplicative regret term $\mathcal{O}(\sqrt{K}/\varepsilon)$ [16] that is often unacceptable for control applications. This naturally leads to the following question: Can a finer trade-off between privacy and regret in RL be achieved, i.e., achieving utility comparable to that of centralized privacy models but without relying on a trusted central server?
Figure 1: Illustration of the operational workflows in the centralized, local, and shuffle privacy models.
Motivated by these, this paper leverages an intermediate trust model of differential privacy, known as shuffle differential privacy (SDP) [18], as a resilient, distributed architecture for private RL. As illustrated in Figure 1, this paradigm uses a secure shuffler to anonymize a batch of locally randomized data before it reaches the central server, and makes sure that the central server can not distinguish any two users’ data. The shuffler is trusted by these users since it can be implemented by trusted cryptographic mixnets or trusted hardware deployed via distributed constructs [19]. This privacy model provides a stronger privacy guarantee than the centralized model but usually suffers a smaller cost than the local model. The key intuition is that additional randomness of the shuffler creates a privacy blanket [20] so that each user can inject much less noise and hide their information in the crowd. A better privacy-utility trade-off has been proven in related fields like (federated) supervised learning [21, 22], and bandits [23, 24]. However, to the best of our knowledge, the application of the shuffle model in RL remains largely unexplored. This gap is due to a fundamental tension between the batch-processing requirement of shuffle privacy and the sequential, online nature of most RL algorithms. Specifically, the shuffle model introduces two compounding challenges. First, by operating on anonymized batches, it imposes an inherent delay, as the agent receives no feedback until an entire batch of interactions is complete. Second, to ensure privacy, the aggregated data is intentionally perturbed with statistical noise. This combination of delayed feedback and noisy observations severely complicates the crucial exploration-exploitation trade-off, posing a significant challenge to developing private RL algorithms that can achieve near-optimal regret while satisfying the SDP constraint.
The batch-based updates required by SDP makes the literature on multi-batch RL highly relevant. This field aims to improve deployment efficiency by minimizing the number of policy switches while maintaining low regret. Approaches in this area are typically either adaptive [25, 26], where batch sizes are determined adaptively based on online data collection, or static [27, 28], where the batch schedule is fixed in advance and offer no intra-batch adaptivity. This line of work has produced algorithms that limit the number of policy switches to an optimal $\mathcal{O}(\log\log K)$, while maintaining an optimal $\tilde{\mathcal{O}}(\sqrt{K})$ regret. The static approach is more relevant to our context, as its core constraint, not using data until a batch is complete, naturally aligns with the shuffle privacy protocol. However, these existing methods are not directly applicable since they can access accurate interaction histories to optimize the policy. In contrast, our algorithm must be robust to the significant noise injected to ensure privacy. This requires a fundamentally different design, one where the batching mechanism and behavior policies are carefully co-designed to control and mitigate the impact of these privacy-induced perturbations.
In this paper, we overcome these challenges by introducing Shuffle Differentially Private Policy Elimination (SDP-PE), the first RL algorithm designed to satisfy the constraint of shuffle privacy. We provide both rigorous theoretical analysis and empirical validation for our method, and the main contributions are threefold:
  • Generic Algorithm Design: Based on the policy elimination idea [27], we develop two key strategies tailored for the shuffle model: a) a staged learning process with exponentially increasing batch sizes. This structure allows for the iterative refinement of an active policy set, and is critical for managing the exploration-exploitation trade-off under the delayed feedback inherent to the shuffle model. b) a “forgetting” mechanism that uses only data from the current stage for model updates. This approach prevents the accumulation of privacy-induced noise across the learning stages.
  • Novel SDP Privatizer for RL: We design a new SDP Privatizer that provides strong utility guarantees for the statistical estimates required in RL. Built upon a shuffle-private binary summation protocol, our privatizer is designed for seamless integration into the SDP-PE framework, ensuring both rigorous privacy and high performance.
[…]
We summarize the comparison of best-known results under three distinct privacy guarantees in Table I: $\varepsilon$-JDP, $\varepsilon$-LDP, and $(\varepsilon,\beta)$-SDP, where $\varepsilon$ and $\beta$ refer to the privacy budget and probability of privacy failure, respectively. Due to space limitations, some proof details are deferred to the appendix.
TABLE I: Results Comparison
Algorithm Privacy model Best-known regret bounds
DP-UCBVI  [29] $\varepsilon$-JDP $\widetilde{\mathcal{O}}\left(\sqrt{XAH^{3}K}+\frac{X^{2}AH^{3}}{\varepsilon}\right)$
$\varepsilon$-LDP $\widetilde{\mathcal{O}}\left(\sqrt{XAH^{3}K}+\frac{X^{2}AH^{3}\sqrt{K}}{\varepsilon}\right)$
SDP-PE (Ours) $(\varepsilon,\beta)$-SDP $\widetilde{\mathcal{O}}\left(\sqrt{XAH^{3}K}+\frac{X^{3}AH^{6}}{\varepsilon}\right)$
The remainder of this paper is organized as follows. In Section II, we provide the preliminaries of episodic RL and adapt Shuffle DP for RL and formulate the problem. Section III presents the designed policy elimination algorithm under SDP constraints. The regret and privacy guarantees are presented in Sections IV and V. The numerical experiment is shown in Section VI. Discussion and conclusion are presented in Section VII and VIII, respectively.
II Preliminaries and Problem formulation
In this section, we review the necessary background on episodic reinforcement learning and shuffle differential privacy, and then formally state the private RL problem.
II-A Episodic Reinforcement Learning
An episodic Markov decision process (MDP) is defined by a tuple $\left(\mathcal{X},\mathcal{A},H,\{P_{h}\}_{h=1}^{H},\{\mathcal{R}_{h}\}_{h=1}^{H},d_{1}\right)$, where $\mathcal{X},\mathcal{A}$ are state and action spaces with respective cardinalities $X$ and $A$, and $H$ is the episode length. At step $h$, the transition function $P_{h}(\cdot|x,a)$ takes a state-action pair and returns a distribution over states, the reward distribution $\mathcal{R}_{h}(x,a)$ is a distribution over $\{0,1\}$ with expectation $r_{h}(x,a)$, and $d_{1}$ is the distribution of initial state. [^2] For simplicity, we assume that the rewards are binary. However, our results naturally extend to $[0,1]$, by replacing our private binary summation mechanism (defined in Section V) with a private summation mechanism for real numbers in $[0,1]$ with similar guarantees. A deterministic policy is defined as a collection $\pi=(\pi_{1},\dots,\pi_{H})$ of policies $\pi_{h}:\mathcal{X}\rightarrow\mathcal{A}$. The value function $V_{h}^{\pi}$ and Q function $Q_{h}^{\pi}$ are defined as: $V_{h}^{\pi}(x)=\mathbb{E}_{\pi}[\sum_{t=h}^{H}r_{t}|x_{h}=x],Q_{h}^{\pi}(x,a)=\mathbb{E}_{\pi}[\sum_{t=h}^{H}r_{t}|x_{h},a_{h}=x,a],\forall x,a\in\mathcal{X}\times\mathcal{A}$. There exists an optimal policy $\pi^{\star}$ such that $V_{h}^{\star}(x)=V_{h}^{\pi^{\star}}(x)=\max_{\pi}V_{h}^{\pi}(x)$ for all $x,h\in\mathcal{X}\times[H]$. [^3] For a positive integer $n$, we define $n]\!:=\!\{1,\ldots,n\}$. The Bellman (optimality) equation follows $\forall h\in[H]:Q_{h}^{\star}(x,a)=r_{h}(x,a)+\max_{a^{\prime}}\mathbb{E}_{x^{\prime}\sim P_{h}(x,a)}[V_{h+1}^{\star}(x^{\prime})].$ The optimal policy is the greedy policy: $\pi^{\star}_{h}(x)=\mathop{\mathrm{argmax}}_{a}Q_{h}^{\star}(x,a),\forall x\in\mathcal{X}.$ For generalization, we define the value function of $\pi$ under MDP transition $p$ and reward function $r^{\prime}$ as $V^{\pi}(r^{\prime},p)$.
We consider an agent (e.g., a controller) tasked with learning to interact with unknown dynamics, modeled as an MDP. The agent interacts with the environment for $K$ episodes. In the security-conscious CPS setting, each episode $k\in[K]$ is modeled as an interaction with a unique system component or user $u_{k}\in\mathcal{U}$, where $\mathcal{U}$ is the user space. For each episode $k\in[K]$, the learning agent determines policy $\pi_{k}$ and sends it to user $u_{k}$ for execution. The output of the execution, a trajectory $S_{k}=\left(x_{h}^{k},a_{h}^{k},r_{h}^{k}\right)_{h\in[H]}$ representing the sensitive data, is sent back to the learning agent for updating the policy. We measure the performance of a learning algorithm by its cumulative regret after $K$ episodes,
$\text{Regret}(K):=\sum_{k=1}^{K}\left[V_{1}^{\star}(x_{1}^{k})-V_{1}^{\pi_{k}}(x_{1}^{k})\right],$ (1)
where $x_{1}^{k}$ is the initial state sampled from $d_{1}$.
II-B Differential Privacy and Shuffle Model
Differential Privacy provides a formal guarantee of data privacy. Consider a general case where $\mathcal{S}$ is the data universe, and we have $n$ unique users. We say $D,D^{\prime}\in\mathcal{S}^{n}$ are neighboring batched datasets if they only differ in one user’s data for some $i\in[n]$. Then, we have the standard definition of DP [5]:
Definition 1 (Differential Privacy (DP))
For $\varepsilon,\beta>0$, a randomized mechanism $\mathcal{M}$ is $(\varepsilon,\beta)$-DP if for all neighboring datasets $D,D^{\prime}$ and any event $E$ in the range of $\mathcal{M}$, we have
$\mathbb{P}[\mathcal{M}(D)\in E]\leq\exp(\varepsilon)\cdot\mathbb{P}[\mathcal{M}(D^{\prime})\in E]+\beta.$
Here, the randomized mechanism $\mathcal{M}$ is an algorithm that takes a dataset as input and uses internal randomness to produce a privatized output, such as a statistic. The special case of $(\varepsilon,0)$-DP is also called pure DP, whereas, for $\beta>0$, $(\varepsilon,\beta)$-DP is referred to as approximate DP.
The shuffle model is an enhanced DP protocol $\mathcal{T}=(\mathcal{E},\mathcal{F},\mathcal{G})$ involving three parts, as shown in Figure 1: (i) a (local) random encoder $\mathcal{E}$ at each user’s side; (ii) a secure shuffler $\mathcal{F}$; and (iii) an analyzer $\mathcal{G}$ at the central server (i.e., the learning agent). In this model, each user $u_{i}$ first locally randomizes their data $D_{i}$ via the encoder $\mathcal{E}$, and sends the resulting messages $\mathcal{E}(D_{i})$ to the shuffler $\mathcal{F}$. The shuffler $\mathcal{F}$ permutes messages from this batch users at random and then reports $\mathcal{F}(\mathcal{E}(D_{1}),\dots,\mathcal{E}(D_{n}))$ to the analyzer. The analyzer $\mathcal{G}$ then aggregates the received messages and outputs desired statistics. In this protocol, the users trust the shuffler but not the analyzer, and the goal is to ensure the output of the shuffler $\mathcal{F}$ on two neighboring datasets is indistinguishable in the analyzer’s view. Here, we define the mechanism $(\mathcal{F}\circ\mathcal{E}^{n})(D)=\mathcal{F}(\mathcal{E}(D_{1}),\dots,\mathcal{E}(D_{n}))$, where $D\in\mathcal{D}^{n}$. Thus, we have the definition of standard shuffle DP [18]:
Definition 2 (Shuffle Differential Privacy (SDP))
A protocol $\mathcal{T}=(\mathcal{E},\mathcal{F},\mathcal{G})$ for $n$ users is $(\varepsilon,\beta)$-SDP if the mechanism $\mathcal{F}\circ\mathcal{E}^{n}$ satisfies $(\varepsilon,\beta)$-DP.
In a sequential data-generation setting, this protocol will be applied across multiple batches. To protect the sequence of outputs across all $B$ batches, we define the (composite) mechanism $\mathcal{M}_{\mathcal{T}}=(\mathcal{F}\circ\mathcal{E}^{L_{1}},\dots,\mathcal{F}\circ\mathcal{E}^{L_{B}}),$ where each mechanism $\mathcal{F}\circ\mathcal{E}^{L_{b}}$ operates on a dataset of $L_{b}$ users. This leads to the following definition $B$-batch SDP [18]:
Definition 3 ($B$-batch SDP)
A $B$-batch shuffle protocol $\mathcal{T}$ is $(\varepsilon,\beta)$-SDP if the mechanism $\mathcal{M}_{\mathcal{T}}$ satisfies $(\varepsilon,\beta)$-DP.
II-C Problem Formulation: Shuffle-Private RL
Here, we formally define the private reinforcement learning problem. We consider an agent interacting with an unknown episodic MDP as described in Section II-A, but under the communication constraint of the $B$-batch shuffle privacy defined in Section II-B.
The core of our problem lies in the information feedback loop. The learning process is partitioned into $B$ batches. In each batch $b$, the agent deploys some policy to a group of $L_{b}$ users, and each user’s resulting interaction trajectory is considered their private data. Crucially, the agent does not observe the raw interaction trajectories. Instead, they are processed by a shuffle protocol, and the agent only receives a privatized, aggregated output for that batch, which it then uses to select policies for the subsequent batch. The central challenge is to design an algorithm that achieves a dual objective:
  • Learning performance: Minimize the cumulative regret $\text{Regret}(K)$.
  • System privacy: Guarantee that the entire sequence of outputs observed by the agent satisfies a $B$-batch $(\varepsilon,\beta)$-SDP guarantee.
This formulation captures the fundamental tension of our work: the agent must learn to make effective sequential decisions despite receiving only delayed, batched, and noisy information.
III Algorithm Design
In this section, we introduce a generic algorithmic framework, Shuffle Differentially Private Policy Elimination (SDP-PE, Algorithm 1). Unlike existing private RL algorithms that rely on the optimism in the face of uncertainty principle [14, 16], SDP-PE builds on the policy elimination idea from [27, 28], which extends action elimination from bandits [30] to RL.
As illustrated in Figure 2, SDP-PE operates in a sequence of stages with exponentially increasing batch sizes, where $L_{b}:=2^{b}$ for $b=1,\dots,B$ and $B=\mathcal{O}(\log K)$. Within each stage, the algorithm maintains a set of “active” policies ($\phi$) and follows a three-step procedure: Crude Exploration, Fine Exploration, and Policy Elimination. The process begins by broadly exploring the environment to build a preliminary model. This model then guides a more refined exploration phase to gather higher-quality data and keep low regret. Finally, using this refined data, it constructs private confidence bounds on the value of each active policy and eliminates any policy that is provably sub-optimal. This cycle of explore-refine-eliminate repeats across stages. As the algorithm proceeds and the batch sizes grow, the confidence intervals shrink, ensuring that only near-optimal policies remain.
The high-level, three-step structure of our algorithm is inspired by the APEVE framework of [27]. However, our objective of minimizing regret under shuffle privacy, rather than minimizing policy switches, necessitates several fundamental innovations in the design as detailed below.
1) Robust estimation via a privatized absorbing MDP. We adapt the technique of using an absorbing MDP from APEVE to handle rarely visited states. The key difference is that in our setting, the set of “infrequently-visited” tuples must be identified from private counts (cf. Assumption 4). This demands a careful construction that is robust to statistical noise, allowing the algorithm to focus its learning on the core, reachable parts of the environment.
2) Optimal trade-offs via exponential batch sizing and “forgetting”. This principle is a core departure from prior work, designed to manage the dual trade-offs of exploration vs. exploitation and privacy vs. regret.
Exploration vs. exploitation: The schedule of increasing batch sizes manages this trade-off. Early, smaller batches prioritize exploration by enabling frequent policy updates to quickly move away from a poor initial policy. As the algorithm’s confidence grows, larger batches are used for longer periods of confident exploitation, leveraging higher-quality data to reliably eliminate sub-optimal policies.
Regret control: Crucially, the exponential growth rate (i.e., $L_{b}\propto 2^{b}$) is mathematically essential. Our analysis shows that this specific schedule ensures two things simultaneously: the non-private regret term matches the lower bound in standard episodic RL, while the privacy cost in regret becomes a logarithmic and additive term. Any other schedule would result in a suboptimal privacy cost.
Forgetting: This batching strategy is paired with a “forgetting” mechanism, using only data from the current stage for estimates, to prevent the accumulation of noise from previous, less accurate stages.
3) A noise-robust fine exploration policy. A third innovation of SDP-PE lies in policy design for the fine exploration phase. Unlike its counterpart in APEVE, our policy is specifically designed to guarantee uniform data coverage over all active policies even in the presence of privacy noise. This is critical for deriving the tighter confidence bounds needed to achieve our final regret.
The following sections will now elaborate on the components of this framework, detailing the private counting mechanisms and each of the three procedural steps.
Figure 2: A visualization of the procedures of SDP-PE (Algorithm 1).
Algorithm 1 Shuffle DP Policy Elimination (SDP-PE)
Parameters: Episode number $K$, universal constant $C$, failure probability $\delta$, privacy budget $\varepsilon>0$ and a Privatizer.
Initialization: $\phi_{1}=\{\text{all deterministic policies}\}$, $\iota=\log(2HAK/\delta)$. Set precision levels $E_{\varepsilon,\delta}$ for Privatizer.
1:for $b=1$ to $B$ do
2:  if $2(\sum_{i=1}^{b}L_{i})\geq K$ then
3:   $L_{b}=\frac{K-2(\sum_{i=1}^{b}L_{i})}{2}$. (o.w. $L_{b}=2^{b}$)
4:  end if
5:  $\mathcal{W}^{b},\!\widetilde{P}^{\text{cru},b},\!\pi_{b}\!=\!\text{CrudeExploration}(\phi_{b},L_{b},\varepsilon,\!\text{Privatizer}).$
6:  $\widetilde{P}^{\text{ref},b},\widetilde{r}^{b}=\text{FineExploration}(\mathcal{W}^{b},\widetilde{P}^{\text{cru},b},\phi_{b},\pi_{b},L_{b},\varepsilon,$ $\text{Privatizer})$
7:  $\psi_{b}=\emptyset$
8:  for $\pi\in\phi_{b}$ do
9:   if $\sup_{\mu\in\phi_{b}}V^{\mu}(\widetilde{r}^{b},\widetilde{P}^{\text{ref},b})-V^{\pi}(\widetilde{r}^{b},\widetilde{P}^{\text{ref},b})\geq 2C\big{(}\sqrt{\frac{XAH^{3}\iota}{L_{b}}}+\frac{X^{3}AH^{5}E_{\varepsilon,\delta}\iota}{L_{b}}\big{)}$ then
10:    Update $\psi_{b}\leftarrow\psi_{b}\cup\{\pi\}$.
11:   end if
12:  end for
13:  $\phi_{b+1}\leftarrow\phi_{b}\backslash\psi_{b}$.
14:end for
III-A Model Estimation via Private Counts in Algorithm 1
SDP-PE employs a model-based approach, relying on a “forgetting” principle to estimate the MDP model in each stage. This means that only the data from the current batch is used for model estimation. While the specific dataset construction varies between the crude exploration and fine exploration steps, as will be detailed later, the general process for estimating the model from a batch trajectories is as follows.
Given a batch dataset $D$ from $n$ users, we define the true visitation counts and cumulative rewards at each step $h$, $N_{h}(x,a,x^{\prime})\!:=\!\sum_{i=1}^{n}\mathds{1}\{x_{h}^{i},a_{h}^{i},x_{h+1}^{i}=x,a,x^{\prime}\}$, similarly $N_{h}(x,a)\!\!:=\!\!\sum_{x^{\prime}\in\mathcal{X}}N_{h}(x,a,x^{\prime})$, and $R_{h}(x,a)\!:=\!\sum_{i=1}^{n}\mathds{1}\{x_{h}^{i},a_{h}^{i}=x,a\}\cdot r_{h}^{i}$. These empirical counts are then processed by a shuffle Privatizer, which then releases their privatized versions, denoted as $\widetilde{N}_{h}(x,a,x^{\prime}),\widetilde{N}_{h}(x,a)$, and $\widetilde{R}_{h}(x,a)$. As formalized in Assumption 4, these private counts are guaranteed to closely approximate the true value, which is essential for reliable learning. The detailed implementation of such Privatizer is justified in Section V.
Assumption 4 (Private counts)
For any privacy budget $\varepsilon>0$ and failure probability $\delta\in(0,1)$, the private counts returned by Privatizer satisfy, for some $E_{\varepsilon,\delta}>0$, with probability at least $1-3\delta$, over all $\left(h,x,a,x^{\prime}\right)$, $|\widetilde{N}_{h}(x,a)\!-\!N_{h}(x,a)|$ $\leq\!\!E_{\varepsilon,\delta},|\widetilde{N}_{h}(x,a,x^{\prime})\!-\!N_{h}(x,a,x^{\prime})|\!\!\leq\!\!E_{\varepsilon,\delta}$, $|\widetilde{R}_{h}(x,a)-R_{h}(x,a)|\leq E_{\varepsilon,\delta}$ and $\widetilde{N}_{h}(x,a)$ $\sum_{x^{\prime}\in\mathcal{X}}$ $\widetilde{N}_{h}(x,a,x^{\prime})$ $\geq$ $N_{h}(x,a)$, $\widetilde{N}_{h}(x,a,x^{\prime})>0$.
Based on these private counts, we define the estimated transition and reward functions for current stage:
$\widetilde{P}_{h}\left(x^{\prime}|x,a\right):=\frac{\widetilde{N}_{h}(x,a,x^{\prime})}{\widetilde{N}_{h}(x,a)},\widetilde{r}_{h}(x,a):=\frac{\widetilde{R}_{h}(x,a)}{\widetilde{N}_{h}(x,a)}.$ (2)
A key property of our Privatizer, enforced by Assumption 4, is that it maintains summation consistency, i.e., $\widetilde{N}_{h}(x,a)=\sum_{x^{\prime}\in\mathcal{X}}\widetilde{N}_{h}(x,a,x^{\prime})$. This construction guarantees that our estimated transition function, $\widetilde{P}_{h}\left(\cdot|x,a\right)$, is always a valid probability distribution.
III-B Crude Exploration in Algorithm 2
Algorithm 2 Crude Exploration
Input: Policy set $\phi$, number of episodes $L$, privacy budget $\varepsilon$ and a Privatizer.
Initialization: $L_{0}=\frac{L}{H},C_{1}=6,\mathcal{W}=\emptyset,\iota=\log(2HAK/\delta)$. $1_{h,x,a}$ is a reward function $r^{\prime}$ where $r^{\prime}_{h^{\prime}}(x^{\prime},a^{\prime})=\mathds{1}\{(h^{\prime},x^{\prime},a^{\prime})=(h,x,a)\}$. $x^{\dagger}$ is an additional absorbing state. $\tilde{P}^{\text{cru}}$ is a transition function over extended space $\mathcal{X}\cup\{x^{\dagger}\}\times\mathcal{A}$, initialized arbitrarily.
Output: Infrequent tuples $\mathcal{W}$, and crude estimated transition function $\tilde{P}^{\text{cru}}$, mixture policy $\pi_{0}$
1:for $h=1$ to $H$ do
2:  Set data set $D^{h}=\emptyset$.
3:  for $(x,a)\in\mathcal{X}\times\mathcal{A}$ do
4:   $\pi^{\text{cru}}_{h,x,a}=\mathop{\mathrm{argmax}}_{\pi\in\phi}V^{\pi}(1_{h,x,a},\tilde{P}^{\text{cru}})$.
5:  end for
6:  Run $\pi^{\text{cru}}_{h}=$ uniform mixture of $\{\pi^{\text{cru}}_{h,x,a}\}_{(x,a)}$ for $L_{0}$ episodes, and add the trajectories into dataset $D^{h}$.
7:  Send batch dataset $D^{h}$ to the Privatizer.
8:  Receive counts $\widetilde{N}^{\text{cru}}_{h}(x,a,x^{\prime})$ and $\widetilde{N}^{\text{cru}}_{h}(x,a)$ for all $(x,a,x^{\prime})$ in $h$-th horizon from Privatizer.
9:  $\mathcal{W}\!=\mathcal{W}\cup\{(h,x,a,x^{\prime})|\widetilde{N}^{\text{cru}}_{h}(x,a,x^{\prime})\!\leq\!C_{1}E_{\varepsilon,\delta}H^{2}\iota\}\!.$
10:  $\tilde{P}^{\text{cru}}=\text{Estimate Transition}(\widetilde{N}^{\text{cru}}_{h},\mathcal{W},x^{\dagger},h,\tilde{P}^{\text{cru}}).$
11:end for
12: Policy $\pi_{0}\leftarrow$ uniform mixture of $\{\pi^{\text{cru}}_{h,x,a}\}_{(h,x,a)}$.
13:return  $\{\mathcal{W},\tilde{P}^{\text{cru}},\pi_{0}\}$.
The first step, crude exploration (Algorithm 2), is designed to build a preliminary, or “crude,” model of the environment $\widetilde{P}^{\text{cru},b}_{h}(x^{\prime}|x,a)$ for any tuple $(h,x,a,x^{\prime})$. The main challenge is that some state-action pairs may be too rarely visited to gather sufficient statistics, especially under privacy constraints. Our strategy is to formally partition the system’s state space by identifying a set of infrequently-visited tuples, denoted as $\mathcal{W}$. The key idea behind is that for the tuples not in $\mathcal{W}$, we can get accurate estimates, and for tuples in $\mathcal{W}$, they have little influence on the value estimate.
To identify this set, we perform a layer-wise exploration. For each step $h$ of episode, we explore every state-action pair $(x,a)$ by constructing a policy, $\pi^{\text{cru},b}_{h,x,a}$, that maximizes the estimated probability of visiting that pair under the current model. We then form a single exploration policy for the layer $\pi^{\text{cru},b}_{h}$, a uniform mixture of these specialized policies. This mixture policy is executed and the resulting trajectories $D^{h}$ are batched and sent to the Privatizer. These returned private counts are then used to empirically identify all tuples visited fewer than $\mathcal{O}(E_{\varepsilon,\delta}H^{2}\iota)$ times, which form the set $\mathcal{W}$.
This set is then used to construct an absorbing MDP, a technique adapted from [27]. As fomally defined below, this method isolates the hard-to-reach states by redirecting their transition probabilities to a single absorbing state, $x^{\dagger}$.
Definition 5 (Abosorbing MDP $P^{\prime}$)
Given $\mathcal{W}$ and $P$, $\forall(h,x,a,x^{\prime})\notin\mathcal{W}$, let $P^{\prime}_{h}(x^{\prime}|x,a)=P_{h}(x^{\prime}|x,a)$, $\forall(h,x,a,x^{\prime})\in\mathcal{W}$, $P^{\prime}_{h}(x^{\prime}|x,a)=0$. For any $(h,x,a)\in[H]\times\mathcal{X}\times\mathcal{A}$,we define $P^{\prime}_{h}(x^{\dagger}|x^{\dagger},a)=1$ and $P^{\prime}_{h}(x^{\dagger}|x,a)=1-\sum_{x^{\prime}\in\mathcal{X}:(h,x,a,x^{\prime})\notin\mathcal{W}}P^{\prime}_{h}(x^{\prime}|x,a).$
The final output is $\widetilde{P}^{\text{cru},b}$ derived from Equation (2), a private and reliable crude estimate of the absorbing MDP. This crude model allows the algorithm to focus its learning capacity on the reachable core of the environment and serves as the foundation for the next step. An auxiliary mixture policy, $\pi_{0}$ is also generated for use in the subsequent phase.
III-C Fine Exploration in Algorithm 3
Algorithm 3 Fine Exploration
Input: Infrequent tuples $\mathcal{W}$, crude estimated transition $\tilde{P}^{\text{cru}}$, policy set $\phi$, auxiliary policy $\pi_{0}$, number of episodes $L$, privacy budget $\varepsilon$ and a Privatizer.
Initialization: $D=\emptyset$. $1_{h,x,a}$ is a reward function $r^{\prime}$ where $r^{\prime}_{h^{\prime}}(x^{\prime},a^{\prime})=\mathds{1}\{(h^{\prime},x^{\prime},a^{\prime})=(h,x,a)\}$. Initialize refined transition estimate $\tilde{P}^{\text{ref}}=\tilde{P}^{\text{cru}}$.
Output: Refined estimated transition function $\tilde{P}^{\text{ref}}$ and reward function $\widetilde{r}$.
1: Construct explorative policy $\pi^{\text{ref}}$ according to Eq. 3.
2: Run $\pi^{\text{ref}}$ for $L$ episodes and $\pi_{0}$ for $L$ episodes, add trajectories to dataset $D$.
3: Send batch dataset $D$ to the Privatizer.
4: Receive private counts $\widetilde{N}^{\text{ref}}_{h}(x,a,x^{\prime})$, $\widetilde{N}^{\text{ref}}_{h}(x,a)$ and $\widetilde{R}^{\text{ref}}_{h}(x,a,x^{\prime})$ for all $(h,x,a,x^{\prime})$ from Privatizer.
5:for $h\in[H]$ do
6:  $\tilde{P}^{\text{ref}}=\text{Estimate Transition}(\widetilde{N}^{\text{ref}}_{h},\mathcal{W},x^{\dagger},h,\tilde{P}^{\text{ref}}).$
[…]
8:end for
9:return  $\tilde{P}^{\text{ref}}$, $\widetilde{r}$.
The second step, fine exploration (Algorithm 3), leverages the crude model $\widetilde{P}^{\text{cru},b}$ to design a more efficient and focused exploration strategy. The key insight is that $\widetilde{P}^{\text{cru},b}$ is reliable enough to construct a uniformly exploration policy, $\pi^{\text{ref},b}$, as defined in Equation (3):
$\pi^{\text{ref},b}=\mathop{\mathrm{argmin}}_{\pi\in\phi_{b}}\sup_{\mu\in\phi_{b}}\sum_{h=1}^{H}\sum_{x,a}\frac{V^{\mu}(1_{h,x,a},\widetilde{P}^{\text{cru},b})}{V^{\pi}(1_{h,x,a},\widetilde{P}^{\text{cru},b})}.$ (3)
This policy is specifically designed to minimize the “worst-case coverage number” over the current set of active policies, which could provide uniform coverage of all remaining policies. To generate data, the users execute both this new policy $\pi^{\text{ref},b}$ and the auxiliary policy $\pi_{0}$ from crude exploration phase. The combined dataset $D$ is then processed by the Privatizer to yield a refined private estimate $\widetilde{P}^{\text{ref},b}$ and reward estimate $\widetilde{r}^{b}$, which are more accurate than their crude counterparts. Notably, different from the APEVE algorithm [27], our fine exploration policy is specifically designed to achieve better policy coverage, leading to tighter confidence bounds and an improved final regret.
The final step is each stage is policy elimination. We evaluate all policies in $\phi_{b}$ using the reliable estimates $\widetilde{P}^{\text{ref},b}$ and $\widetilde{r}^{b}$, and update the active policy set by eliminating all policies whose upper confidence bound (UCB) is less than the lower confidence bound (LCB) of any other active policy. The confidence intervals shrink along with the policy elimination step, ensuring that the optimal policy stays in the active policy set with high probability, which guarantees that the algorithm eventually converges to the optimal policy.
IV Regret Guarantee
The main theoretical result of this paper is the following regret bound for our SDP-PE algorithm.
Theorem 6 (Regret bound of SDP-PE)
For any privacy budget $\varepsilon>0$ and failure probability $\delta\in(0,1)$, and any Privatizer that satisfies Assumptions 4, with probability at least $1-9\delta$, the regret of SDP-PE (Algorithm 1) is
$\text{Regret}(K)\leq\tilde{\mathcal{O}}\left(\sqrt{XAH^{3}K}+X^{3}AH^{5}E_{\varepsilon,\delta}\right).$
The core of the regret analysis is to construct a uniform policy evaluation bound that holds for all active policies at each stage of the algorithm. In each stage $b$, the active policy set is $\phi_{b}$. Assume that for any $\pi\in\phi_{b}$, the value $V^{\pi}(r,P)$ can be estimated up to an error $\xi_{b}$ by $V^{\pi}(\widetilde{r}^{b},\widetilde{P}^{\text{ref},b})$, then if we eliminate all policies that are at least $2\xi_{b}$ sub-optimal in the sense of $\max_{\pi\in\phi_{b}}V^{\pi}(\widetilde{r}^{b},\widetilde{P}^{\text{ref},b})$, the optimal policy will never be eliminated and all remaining policies will be at most $4\xi_{b}$ sub-optimal. Summing the regret across all stages yields the total regret:
$\text{Regret}(K)\leq 2HL_{1}+\sum_{b=2}^{B}2L_{b}\cdot 4\xi_{b-1}.$ (4)
The following lemma gives an upper bound of $\xi_{b}$ using our private estimate $\widetilde{P}^{\text{ref},b}$ and $\widetilde{r}^{b}$ of the absorbing MDP.
Lemma 7
With probability $1-9\delta$, it holds that for any stage $b$ and $\pi\in\phi_{b}$,
$|V^{\pi}\!(r,\!P)-V^{\pi}\!(\widetilde{r}^{b}\!,\widetilde{P}^{\text{ref},b})|\!\leq\!\tilde{\mathcal{O}}\bigg{(}\!\sqrt{\frac{XAH^{3}}{L_{b}}}+\frac{X^{3}AH^{5}E_{\varepsilon,\delta}}{L_{b}}\!\!\bigg{)}\!.$
By substituting $\xi_{b}$ in Equation 4, we will have
$\text{Regret}(K)\leq\tilde{\mathcal{O}}\left(\sum_{b=2}^{B}L_{b}\cdot\left(\sqrt{\frac{XAH^{3}}{L_{b-1}}}+\frac{X^{3}AH^{5}E_{\varepsilon,\delta}}{L_{b-1}}\right)\right).$ (5)
Here, the carefully designed exponential batch schedule becomes essential. The total regret is a sum over the stages, composed of a non-private term proportional to $L_{b}/\sqrt{L_{b-1}}$ and a private cost term proportional to $L_{b}/L_{b-1}$. By choosing an exponential schedule where $L_{b}/\sqrt{L_{b-1}}$ forms a geometric series and $L_{b}/L_{b-1}$ is a constant, we get a great balance between privacy cost and non-private regret, i.e., an optimal non-private term and an additive private cost term on regret bound in Theorem 6. Any other schedule, such as a fixed or polynomially increasing batch size, would fail to balance both components, resulting in a suboptimal regret dominated by either the non-private or the private cost.
The proof reduces to proving Lemma 7. We leverage the absorbing MDP $P^{\prime}$ as the key intermediate bridge, and then decompose this total error into three components, “Model Bias” $|V^{\pi}(r,P)-V^{\pi}(r,P^{\prime})|$, “Reward Error” $|V^{\pi}(r,P^{\prime})-V^{\pi}(\widetilde{r}^{b},P^{\prime})|$, and “Model Variance” $|V^{\pi}(\widetilde{r}^{b},P^{\prime})-V^{\pi}(\widetilde{r}^{b},\widetilde{P}^{\text{ref},b})|$. We will now sketch the bounds for “Model Bias” and “Model Variance”. The Reward Error can be bounded as a lower-order term using a similar analysis.
IV-A “Model Bias”: Difference between $P$ and $P^{\prime}$
To analyze the difference between the true MDP with $P$ and the absorbing MDP with $P^{\prime}$, we rely on the the properties of the crude transition estimate $\widetilde{P}^{\text{cru},b}$, and then reduce the problem to proving it difficult to visit those infrequently visited tuples.
Property of $\widetilde{P}^{\text{cru},b}$. In $b$-th stage, if the private visitation count $\widetilde{N}^{\text{cru},b}_{h}(x,a,x^{\prime})$ for a tuple $(h,x,a,x^{\prime})$ exceeds $\mathcal{O}(E_{\varepsilon,\delta}H^{2}\iota)$, the following approximation result holds with high probability, $(1-\frac{1}{H})\cdot\widetilde{P}^{\text{cru},b}_{h}(x^{\prime}|x,a)\leq P^{\prime}_{h}(x^{\prime}|x,a)\leq(1+\frac{1}{H})\cdot\widetilde{P}^{\text{cru},b}_{h}(x^{\prime}|x,a)$, which can be proven using Bernstein’s inequality. By the construction of $\mathcal{W}$, and $P^{\prime},\widetilde{P}^{\text{cru},b}$, the above equation holds for any $(h,x,a,x^{\prime})$. Consequently, for any $(h,x,a)\in[H]\times\mathcal{X}\times\mathcal{A},\pi\in\phi_{b}$, we have
$\frac{1}{4}V^{\pi}(1_{h,x,a},\widetilde{P}^{\text{cru},b})\leq V^{\pi}(1_{h,x,a},P^{\prime})\leq 3V^{\pi}(1_{h,x,a},\widetilde{P}^{\text{cru},b}).$
Uniform bound on $|V^{\pi}(r,P)-V^{\pi}(r,P^{\prime})|$. Next, we aim to bound $\sup_{\pi\in\phi_{b}}\sup_{r^{\prime}}|V^{\pi}(r^{\prime},P)-V^{\pi}(r^{\prime},P^{\prime})|$. This leads to bounding $\sup_{\pi\in\phi_{b}}\mathbb{P}_{\pi}[\mathcal{B}]$, where the event $\mathcal{B}$ occurs when the trajectory visits any infrequently visited tuples in $\mathcal{W}$, which describes the main difference between absorbing MDP and the true MDP. By the definition of $\mathcal{W}$, we can show that these tuples are difficult to visit for any policy in $\phi_{b}$, i.e., with high probability, $\sup_{\pi\in\phi_{b}}\mathbb{P}_{\pi}[\mathcal{B}]\leq\tilde{\mathcal{O}}\left(\frac{X^{3}AH^{4}E_{\varepsilon,\delta}}{L_{b}}\right)$. Using this observation, we bound the bias with proof in Appendix Proof of Lemma 8.
Lemma 8
With high probability, for any policy $\pi\in\phi_{b}$ and reward function $r^{\prime}$, it holds that
$0\leq V^{\pi}(r^{\prime},P)-V^{\pi}(r^{\prime},P^{\prime})\leq\widetilde{\mathcal{O}}\left(\frac{X^{3}AH^{5}E_{\varepsilon,\delta}}{L_{b}}\right).$
IV-B “Model Variance”: Difference between $P^{\prime}$ and $\widetilde{P}^{\text{ref},b}$
This term bounds how much our refined estimate, $\widetilde{P}^{\text{ref},b}$, deviates from the absorbing MDP, $P^{\prime}$. Using the standard simulation lemma from [31], this variance term can be roughly bounded by $\widetilde{\mathcal{O}}\left(\sqrt{\sum_{h,x,a}\frac{V^{\pi}(1_{h,x,a},P^{\prime})}{N_{h}^{\text{ref},b}(x,a)}}\right)$. Meanwhile, the visitation counts is roughly related to the exploration policy, $N_{h}^{\text{ref},b}(x,a)\propto V^{\pi^{\text{ref},b}}(1_{h,x,a},P^{\prime})$. Thus, the key to bounding this term is our fine exploration policy, $\pi^{\text{ref},b}$. By its design, this policy minimizes the worst-case coverage number $\max_{\pi\in\phi_{b}}\sum_{h,x,a}\frac{V^{\pi}(1_{h,x,a},P^{\prime})}{V^{\pi^{\text{ref},b}}(1_{h,x,a},P^{\prime})}$ over all active policies, with a order of $\mathcal{O}(XAH)$. This ensures sufficient exploration of all relevant state-action pairs, leading to the following guarantee (proof in Appendix Proof of Lemma 9).
Lemma 9
With high probability, for any policy $\pi\in\phi_{b}$ and reward function $r^{\prime}$, it holds that
$|V^{\pi}\!(r^{\prime}\!,\!P^{\prime})-V^{\pi}\!(r^{\prime}\!,\!\widetilde{P}^{\text{ref},b})|\!\leq\!\widetilde{\mathcal{O}}\!\bigg{(}\!\sqrt{\frac{XAH^{3}}{L_{b}}}+\frac{X^{2}AH^{2}E_{\varepsilon,\delta}}{L_{b}}\!\bigg{)}\!.$
V Privacy Guarantee
In this section, we present a secure communication protocol, a shuffle Privatizer adapted from the shuffle binary summation mechanism introduced by [18]. We first describe the mechanism satisfying the Assumption 4, and then show that Algorithm 1, when instantiated with this Privatizer, satisfies the $(\varepsilon,\beta)$-SDP guarantee.
To secure the data of the batch users ($D^{h}$ in Crude Exploration and $D$ in Fine Exploration), we privatize the visitation counts and rewards. Given $n$ users in a batch, we outline the process for a single counter $\widetilde{N}_{h}(x,a,x^{\prime})$ of a specific $(h,x,a,x^{\prime})$. The procedure for other counts like $\widetilde{N}_{h}(x,a)$ and rewards $\widetilde{R}_{h}(x,a)$ is analogous.
Firstly, we allocate privacy budget $\varepsilon^{\prime}=\frac{\varepsilon}{3H}$ and failure probability $\beta^{\prime}=\frac{\beta}{HXA}>0$ to one counter, and define a threshold $\tau=\mathcal{O}(\log(1/\beta^{\prime})/\varepsilon^{\prime 2})$ that controls noise addition. On the local side, each user encodes their data using a randomizer $\mathcal{E}$: if $n\leq\tau$, user $i$ encodes local count as $z_{i}=\mathds{1}_{h}\left(x,a,x^{\prime}\right)+\sum_{j=1}^{m}y_{j}$, where $\{y_{j}\}_{j=1}^{m}$ are i.i.d. sampled, $y_{j}\sim\text{Bernoulli}(1/2)$, and $m=\lceil\frac{\tau}{n}\rceil$; otherwise, the local output is $z_{i}=\mathds{1}_{h}\left(x,a,x^{\prime}\right)+y$ where $y\sim\text{Bernoulli}(\frac{\tau}{2n})$. The user then sends its local message $z_{i}$ to a secure shuffler $\mathcal{F}$, which permutes their messages randomly and then forwards them to the analyzer $\mathcal{G}$.
On the analyzer side, it performs aggregation and post-process steps. It firstly aggregates the shuffled messages into an initial noisy count, denoted as $\ddot{N}_{h}(x,a,x^{\prime})$. If $n\leq\tau$, $\ddot{N}_{h}(x,a,x^{\prime})=\sum_{i=1}^{n}z_{i}-\lceil\frac{\tau}{n}\rceil\cdot\frac{n}{2}$, otherwise, $\ddot{N}_{h}(x,a,x^{\prime})=\sum_{i=1}^{n}z_{i}-\frac{\tau}{2}$. The noisy visitation counts are then post-processed to ensure they satisfy Assumption 4, and generate the final counts $\widetilde{N}_{h}(x,a,x^{\prime})$. The private cumulative rewards $\widetilde{R}_{h}(x,a)$ are directly obtained from the aggregation output.
For the post-processing steps for $\ddot{N}_{h}$, we will adapt the techniques from [15, 29]. Firstly, we solve the linear optimization problem for all $\left(x,a\right)$ below.
$\min t\,\,\,\,\text{s.t.}$ $\,\,n(x^{\prime})\geq 0,|n(x^{\prime})-\ddot{N}_{h}(x,a,x^{\prime})|\leq t,\forall x^{\prime},$ (6)
$|\sum\nolimits_{x^{\prime}\in\mathcal{X}}n(x^{\prime})-\ddot{N}_{h}(x,a)|\leq\frac{E_{\varepsilon,\delta}}{4}.$
Let $\bar{N}_{h}(x,a,x^{\prime})$ denote a minimizer of this problem, we define $\bar{N}_{h}(x,a)=\sum_{x^{\prime}\in\mathcal{X}}\bar{N}_{h}(x,a,x^{\prime})$. By adding some term, as done below, the private counts $\widetilde{N}_{h}(x,a)$ never underestimate the respective true counts:
$\widetilde{N}_{h}\!(x,\!a)\!=\!\bar{N}_{h}\!(x,\!a)\!+\!\frac{E_{\varepsilon,\delta}}{2},\widetilde{N}_{h}\!(x,\!a,\!x^{\prime})\!=\!\bar{N}_{h}\!(x,\!a,\!x^{\prime})\!+\!\frac{E_{\varepsilon,\delta}}{2X}.$
Suppose $\ddot{N}_{h}(x,a)$ satisfies $|\ddot{N}_{h}(x,a,x^{\prime})-N_{h}(x,a,x^{\prime})|\leq\frac{E_{\varepsilon,\delta}}{4}$, $|\ddot{N}_{h}(x,a)-N_{h}(x,a)|\leq\frac{E_{\varepsilon,\delta}}{4}$, for all $\left(h,x,a,x^{\prime}\right)$, with probability $1-2\delta$. Then, the derived $\widetilde{N}_{h}(x,a),\widetilde{N}_{h}(x,a,x^{\prime})$ satisfy Assumption 4.
The result of our shuffle Privatizer is a set of reliable statistics, with formal guarantees summarized in the following Lemma with proof in Appendix Proof of Lemma 10.
Lemma 10 (Shuffle DP Privatizer)
For any $\varepsilon\!\in\!(0,1)$, $\beta\!\in\!(0,1)$, the Shuffling Privatizer satisfies $(\varepsilon,\beta)$-SDP and Assumption 4 with $E_{\varepsilon,\delta}=\tilde{\mathcal{O}}\left(\frac{H}{\varepsilon}\right)$.
As corollaries of Theorem 6, we obtain the regret and privacy guarantees for SDP-PE using the shuffle Privatizer.
Theorem 11 (Regret Bound under SDP)
For any $\varepsilon\in(0,1),\beta\in(0,1)$, under the Shuffling Privatizer, SDP-PE satisfies $(\varepsilon,\beta)$-SDP. Furthermore, we obtain $\text{Regret}(K)\leq\tilde{\mathcal{O}}\left(\sqrt{XAH^{3}K}+\frac{X^{3}AH^{6}}{\varepsilon}\right)$ with high probability.
VI Numerical Experiment
In this section, we conduct simple numerical experiments to verify our theoretical results for SDP-PE (Algorithm 1).
VI-A Settings
Due to the algorithm’s computational complexity, the simulation is conducted on a simplified “RiverSwim” [32] environment, a standard benchmark with challenging exploration requirements, illustrated in Figure 3. The system consists of four states and two actions “left” and “right”, i.e., $X=4$ and $A=2$. In each episode, the agent starts from the leftmost state and executes its policy. At each step of episode $k$, if the agent chooses the action “left”, it will always succeed. Otherwise, it may fail. It only receives a small amount of reward if it reaches the leftmost state while obtaining a much larger reward if it arrives at the rightmost state. Thus, the MDP naturally requires sufficient exploration to obtain the optimal policy. Each episode is reset every $H=6$ steps.
Figure 3: Simplified RiverSwim MDP – solid and dotted arrows denote the transitions under actions “right” and “left”, respectively.
We evaluate the performance of SDP-PE and compare it against the state-of-the-art private algorithm, private UCB-VI under JDP and LDP constraints [29], which updates the policy in every episode, as well as their corresponding non-private versions (PE and UCB-VI). We conduct 20 independent experiments under two privacy regimes $\varepsilon=0.1$ and $\varepsilon=1$, each running $K=2\cdot 10^{4}$ episodes.
VI-B Results and Analysis
Figure 4: Cumulative regret vs. Episode under private RL algorithms
We plot the cumulative regret versus the number of episodes, along with the standard deviation, in Figure 4. Our experimental results offer several key insights that align with our theoretical results.
First, in the non-private setting, UCB-VI, as a theoretically optimal method with per-episode updates, unsurprisingly demonstrates the best performance. Our non-private PE algorithm achieves comparable results, confirming that an effective policy can be learned even with a less frequent, i.e., logarithmic update schedule, and a data-forgetting mechanism.
Second, under both private regimes $(\varepsilon=0.1$ and $\varepsilon=1)$, SDP-PE consistently achieves substantially lower cumulative regret than UCBVI-LDP. This empirically verifies our primary goal of attaining a superior performance-privacy trade-off compared to the local privacy model.
Finally, the experiments also highlight a nuanced trade-off within the regret bounds. While the dominant term of SDP-PE’s regret matches the non-private lower bound and the optimal JDP rate, a gap exists due to the privacy-dependent term’s higher dependency on model parameters $(X,A,H)$. This is visible in the plots, where SDP-PE is slightly higher than that of UCBVI-JDP. This gap reflects the cost of reduced data utilization in our batch-based approach and the absence of a trusted central server.
VII Discussion
Here, we discuss the broader implications of our results in the context of the private learning for the CPS literature.
Performance-privacy trade-off. Our SDP-PE algorithm establishes a favorable regret-privacy trade-off, with a regret bound whose leading term matches the non-private lower bound, $\Omega(\sqrt{XAH^{3}K})$. Crucially, our algorithm significantly outperforms the optimal LDP result [29] by making the privacy costs’ dependency on $1/\varepsilon$ additive –rather than multiplicative– with respect to $K$. The utility matches that of the JDP case, but is achieved within a more resilient, distributed architecture that lacks a single point of failure–a critical feature for robust CPS.
The shuffle model as a resilient privacy architecture. The shuffle model serves as an intermediate trust model between the centralized model (JDP) and the local model (LDP) for privacy-preserving learning in CPS. Firstly, a key theoretical insight is that any RL algorithm satisfying SDP also satisfies JDP via Billboard lemma in [33]. Therefore, the JDP lower bound of $\widetilde{\Omega}(\sqrt{XAH^{3}K}+\frac{XAH}{\varepsilon})$ established by [14], also applies to our setting. It remains an open question whether the lower bound is tight. Secondly, while our protocol involves local noise injection, it does not satisfy formal LDP. Meanwhile, due to the “privacy blanket”, the privacy preservation becomes less costly as the batch size grows and local noise decreases in the later stages.
Computational efficiency for CPS deployment. A key consideration for deploying any control algorithm in a real-world CPS is its computational feasibility. A limitation of SDP-PE is its current computation complexity, as it requires explicitly evaluating a potentially large set of active policies, and solving constrained optimization problems to construct exploration policies in each stage. Future work could improve efficiency by designing, e.g., a constrained policy search module inspired by [28] that implicitly optimize over a constrained policy space rather than maintaining an explicit set.
VIII Conclusion and Future Work
In this paper, we introduced Shuffle Differentially Private Policy Elimination (SDP-PE), the first generic algorithm for reinforcement learning under the shuffle privacy model. Our method adapts the policy elimination paradigm to the private learning setting by introducing a novel exponential batching schedule paired with a ”forgetting” mechanism. When instantiated with our proposed Privatizer, SDP-PE achieves a regret of $\widetilde{\mathcal{O}}\left(\sqrt{XAH^{3}K}+\frac{X^{3}AH^{6}}{\varepsilon}\right)$, matching the information-theoretic lower bound up to lower-order terms. Notably, this result achieves utility comparable to the centralized (JDP) model but under a stronger privacy guarantee, while significantly outperforming the local (LDP) model. Our numerical experiments also corroborate these findings and demonstrate the practical advantages of our method. This work thus demonstrates the viability of the shuffle model for achieving a favorable privacy-regret trade-off in data-driven control for advanced CPS. This work opens several avenues for future research. Key directions include improving the computational efficiency of the algorithm for real-time CPS deployment, extending our theoretical guarantees to the setting of function approximation, and exploring the application of alternative privacy notions, such as Rényi DP.
Supporting Lemmas
We first give several supporting lemmas to help analysis.
Lemma 12 (Lemma F.4 in [31])
Let $F_{i}$ for $i=1\cdots$ be a filtration and $X_{1},\cdots,X_{n}$ be a sequence of Bernoulli random variables with $\mathbb{P}\left(X_{i}=1\mid F_{i-1}\right)=P_{i}$ with $P_{i}$ being $F_{i-1}$-measurable and $X_{i}$ being $F_{i}$ measurable. It holds that $\mathbb{P}\left[\exists n:\sum_{t=1}^{n}X_{t}<\sum_{t=1}^{n}P_{t}/2-W\right]\leq e^{-W}.$
Lemma 13 (Lemma 1 in [28])
Let $d>0$ be an integer. Let $\mathcal{X}\subset(\Delta^{d})^{m}$. Then there exists a distribution $\mathcal{D}$ over $\mathcal{X}$, such that $\max_{x=\{x_{i}\}_{i=1}^{dm}\in\mathcal{X}}\sum_{i=1}^{dm}\frac{x_{i}}{y_{i}}=md,$ where $y=\left\{y_{i}\right\}_{i=1}^{dm}=\mathbb{E}_{x\sim\mathcal{D}}[x]$. Moreover, if $\mathcal{X}$ has a boundary set $\partial\mathcal{X}$ with finite cardinality, we can find an approximation solution for $\mathcal{D}$ in poly $(|\partial\mathcal{X}|)$ time.
Proof of Lemma 8
This section provides a detailed proof for bounding the ”Model Bias”—the value difference between the true MDP ($P$) and our absorbing MDP ($P^{\prime}$). The proof proceeds in three main steps. First, we show that our private crude model ($\widetilde{P}^{\text{cru},b}$) is multiplicatively close to the absorbing model for all well-explored tuples. Second, we leverage this to show that for any active policy, the probability of visiting an ”infrequently-visited” tuple is small. Finally, we use this low probability to bound the total difference in value.
We begin by establishing the multiplicative concentration of our private model estimate.
Lemma 14
With probability at least $1-4\delta$, for all $b$, $\forall(h,x,a,x^{\prime})$ such that $(h,x,a,x^{\prime})\notin\mathcal{W}$, it holds that, $(1-\frac{1}{H})\cdot\widetilde{P}^{\text{cru},b}_{h}(x^{\prime}|x,a)\leq P^{\prime}_{h}(x^{\prime}|x,a)\leq(1+\frac{1}{H})\cdot\widetilde{P}^{\text{cru},b}_{h}(x^{\prime}|x,a).$ And $\forall(h,x,a,x^{\prime})\in\mathcal{W}$, $P^{\prime}_{h}(x^{\prime}|x,a)=\widetilde{P}^{\text{cru},b}_{h}(x^{\prime}|x,a)=0$.
Proof:
As the algorithm design, $\widetilde{N}_{h}^{\text{cru},b}(x,a,x^{\prime})\geq C_{1}E_{\varepsilon,\delta}H^{2}\iota$ and $\widetilde{P}^{\text{cru},b}_{h}\left(x^{\prime}|x,a\right)=\frac{\widetilde{N}_{h}^{\text{cru},b}(x,a,x^{\prime})}{\widetilde{N}_{h}^{\text{cru},b}(x,a)}$ for all $(h,x,a,x^{\prime})\notin\mathcal{W}$, then
$\left|P^{\prime}_{h}\left(x^{\prime}|x,a\right)-\widetilde{P}^{\text{cru},b}_{h}\left(x^{\prime}|x,a\right)\right|$
$\leq\sqrt{\frac{2\widetilde{P}^{\text{cru},b}_{h}\left(x^{\prime}|x,a\right)\iota}{\widetilde{N}_{h}^{\text{cru},b}(x,a)}}+\frac{4E_{\varepsilon,\delta}\iota}{\widetilde{N}_{h}^{\text{cru},b}(x,a)}$
$\leq\!\bigg{(}\sqrt{\frac{2\iota}{\widetilde{N}_{h}^{\text{cru},b}(x,a,x^{\prime})}}\!+\!\frac{4E_{\varepsilon,\delta}\iota}{\widetilde{N}_{h}^{\text{cru},b}(x,a,x^{\prime})}\!\bigg{)}\!\cdot\!\widetilde{P}^{\text{cru},b}_{h}\!\left(x^{\prime}|x,a\right)$
$\leq\frac{1}{H}\cdot\widetilde{P}^{\text{cru},b}_{h}\left(x^{\prime}|x,a\right).$
The first line is derived from the Bernstein inequality, and the third line is due to the definition of $\mathcal{W}$ and the choice of $C_{1}=6$. The proof is the same for the left side of the inequality. ∎
Next, we show that the multiplicative closeness of the transition models extends to the value functions computed under them. This is a key step that allows us to reason about policy performance using our crude estimate.
Lemma 15
Conditioned on the event in Lemma 14, for any policy $\pi$ and any $(h,x,a)\in[H]\times\mathcal{X}\times\mathcal{A}$, it holds that
$\frac{1}{4}V^{\pi}(1_{h,x,a},\widetilde{P}^{\text{cru},b})\leq V^{\pi}(1_{h,x,a},P^{\prime})\leq 3V^{\pi}(1_{h,x,a},\widetilde{P}^{\text{cru},b}).$
Proof:
Under the absorbing MDP, for any trajectory $S=\{x_{1},a_{1},\dots,x_{h},a_{h}\}$ truncated at time step $h$ such that $(x_{h},a_{a})\in\mathcal{W}$, we have $x_{h^{\prime}}\neq x^{\dagger}$ for any $h^{\prime}\leq h-1$. It holds that
$\mathbb{P}[S|P^{\prime},\pi]=\prod_{j=1}^{h}\pi_{j}(a_{j}|x_{j})\times\prod_{j=1}^{h-1}P^{\prime}_{j}(x_{j+1}|x_{j},a_{j})$ (7)
$\leq\left(1+\frac{1}{H}\right)^{H}\prod_{j=1}^{h}\pi_{j}(a_{j}|x_{j})\times\prod_{j=1}^{h-1}\widetilde{P}^{\text{cru},b}_{j}(x_{j+1}|x_{j},a_{j})$
$\leq 3\mathbb{P}[S|\widetilde{P}^{\text{cru},b},\pi].$
The second line is due to Lemma 14. Let $\mathcal{S}_{h,x,a}$ be the set of truncated trajectories such that $(x_{h},a_{h})=(x,a)$. Then
$V^{\pi}(1_{h,x,a},P^{\prime})=\sum_{S\in\mathcal{S}_{h,x,a}}\mathbb{P}[S|P^{\prime},\pi]$ (8)
$\leq 3\sum_{S\in\mathcal{S}_{h,x,a}}\mathbb{P}[S|\widetilde{P}^{\text{cru},b},\pi]=3V^{\pi}(1_{h,x,a},\widetilde{P}^{\text{cru},b}).$
The left side of the lemma can be proven in a similar way, with $\left(1-\frac{1}{H}\right)^{H}\geq 1/4$ when $H\geq 2$. ∎
Follow the proof sketch, With these properties established, we can now bound the probability of a trajectory encountering any infrequently-visited tuple (the event $\mathcal{B}$). This is the crucial step in bounding the model bias.
Lemma 16
Conditioned on the event in Lemma 14, with high probability at least $1-7\delta$, for any stage $b\in[B]$, $\sup_{\pi\in\phi_{b}}\mathbb{P}_{\pi}[\mathcal{B}]\leq\mathcal{O}\left(\frac{E_{\varepsilon,\delta}X^{3}AH^{4}\iota}{L_{b}}\right)$.
Proof:
Recall that the event $\mathcal{B}$ is that the trajectory falls some infrequently visited tuple $(h,x,a,x^{\prime})$ in $\mathcal{W}$, We can decompose $\mathcal{B}$ into a disjoint union of $\{\mathcal{B}_{h}\}_{h=1}^{H}$, where $\mathcal{B}_{h}$ is the event that $(h,x_{h},a_{h},x_{h+1})\in\mathcal{W}$ and $(h^{\prime},x_{h^{\prime}},a_{h^{\prime}},x_{h^{\prime}+1})\notin\mathcal{W}$ for all $h^{\prime}\leq h-1$; and $\mathbb{P}(\mathcal{B})=\sum_{h=1}^{H}\mathbb{P}(\mathcal{B}_{h})$.
A key first step is to show that the probability of this event is the same under the true MDP ($P$) and the absorbing MDP ($P^{\prime}$). This is because the two models only differ after an infrequently-visited tuple is encountered.
$\mathbb{P}[\mathcal{B}_{h}|P,\pi]=\sum_{S\in\mathcal{B}_{h}}\mathbb{P}[S|P,\pi]$
$=\sum_{S_{:h+1}\in\mathcal{B}_{h}}\mathbb{P}[(x_{1},a_{1},\dots,x_{h},a_{h})|P,\pi]P_{h}(x_{h+1}|x_{h},a_{h})$
$=\sum_{S_{:h+1}\in\mathcal{B}_{h}}\mathbb{P}[(x_{1},a_{1},\dots,x_{h},a_{h})|P^{\prime},\pi]P_{h}(x_{h+1}|x_{h},a_{h})$
$=\sum_{x\in\mathcal{X},a}V^{\pi}(1_{h,x,a},P^{\prime})\sum_{x^{\prime}\in\mathcal{X}:(h,x,a,x^{\prime})\in\mathcal{W}}P_{h}(x^{\prime}|x,a)$
$=\sum_{x\in\mathcal{X},a}V^{\pi}(1_{h,x,a},P^{\prime})P^{\prime}_{h}(x^{\dagger}|x,a)$
$=\mathbb{P}[\mathcal{B}_{h}|P^{\prime},\pi].$
In the first line, $S_{:h+1}$ means the trajectory $S$ truncated at $x_{h+1}$. The second line is because for $(h,x,a,x^{\prime})\notin\mathcal{W}$, $P=P^{\prime}$. The third line is because there is a bijection between trajectories that arrive at $(h,x,a)$ under absorbing MDP and trajectories in $\mathcal{B}_{h}$ that arrive at the same tuple under the original MDP. The last equations follow the definition of $P^{\prime}$ and $\mathcal{B}_{h}$.
The analysis then reduces to bounding $\mathbb{P}[\mathcal{B}_{h}|P^{\prime},\pi]$. We can express this as the probability of visiting any state $x$ at step $h$, multiplied by the conditional probability of transitioning to an infrequent tuple from there.
$\sup_{\pi\in\phi_{b}}\mathbb{P}[\mathcal{B}_{h}|P,\pi]=\sup_{\pi\in\phi_{b}}\mathbb{P}[\mathcal{B}_{h}|P^{\prime},\pi]$ (9)
$=\sup_{\pi\in\phi_{b}}\sum_{x\in\mathcal{X}}V^{\pi}(1_{h,x},P^{\prime})\max_{a\in\mathcal{A}}P^{\prime}_{h}(x^{\dagger}|x,a)$
$=\sup_{\pi\in\phi_{b}}\sum_{x\in\mathcal{X}}V^{\pi}(1_{h,x},P^{\prime})\max_{a\in\mathcal{A}}\mathbb{P}[\mathcal{W}|(h,x,a)],$
where $\mathbb{P}[\mathcal{W}|(h,x,a)]$ represents the conditional probability of entering $\mathcal{W}$ at time step $h$ via $(x_{h},a_{h})=(x,a)$.
Firstly, we deal with the state visiting issue. If the expected number of visits to a state is low, its contribution to the total probability is small. Since $\pi^{cru}_{h,x,a}=\mathop{\mathrm{argmax}}_{\pi\in\phi_{b}}V^{\pi}(1_{h,x,a},\widetilde{P}^{\text{cru},b})$, and using Lemma 15, then
$V^{\pi^{cru}_{h,x,a}}(1_{h,x,a},P^{\prime})$ $\geq\frac{1}{4}V^{\pi^{cru}_{h,x,a}}(1_{h,x,a},\widetilde{P}^{\text{cru},b})$ (10)
$\geq\frac{1}{12}\sup_{\pi\in\phi_{b}}V^{\pi}(1_{h,x,a},P^{\prime}).$
Since $\pi^{cru}_{h}$ chooses each $\pi^{cru}_{h,x,a}$ with probability $\frac{1}{XA}$ for any $(x,a)\in\mathcal{X}\times\mathcal{A}$, we have,
$V^{\pi^{cru}_{h}}(1_{h,x,a},P^{\prime})$ $\geq\frac{1}{12XA}\sup_{\pi\in\phi_{b}}V^{\pi}(1_{h,x},P^{\prime}).$ (11)
The equality is because we only select among deterministic policies. By Lemma 12 and a union bound, we have with probability $1-\delta$, for any $(x,a)\in\mathcal{X}\times\mathcal{A}$,
$N^{\text{cru},b}_{h}(x,a)$ $\geq\frac{L_{b}}{H}\cdot\frac{V^{\pi^{cru}_{h}}(1_{h,x,a},P)}{2}-\iota$ (12)
$\geq\frac{L_{b}\cdot\text{sup}_{\pi\in\phi_{b}}V^{\pi}(1_{h,x},P^{\prime})}{24HXA}-\iota.$
For fixed $(x,a)$, if $\frac{L_{b}\cdot\text{sup}_{\pi\in\phi_{b}}V^{\pi}(1_{h,x},P^{\prime})}{24HXA}\leq 2\iota$, we have that
$\text{sup}_{\pi\in\phi_{b}}V^{\pi}(1_{h,x},P^{\prime})\leq\frac{48HXA\iota}{L_{b}}.$ (13)
As a result, we have the bound for any $\pi\in\phi_{b}$,
$V^{\pi}(1_{h,x},P^{\prime})\max_{a\in\mathcal{A}}\mathbb{P}[\mathcal{W}|(h,x,a)]\leq V^{\pi}(1_{h,x},P^{\prime})$ (14)
$\leq\sup_{\pi\in\phi_{b}}V^{\pi}(1_{h,x},P^{\prime})\leq\frac{48HXA\iota}{L_{b}}.$
Then we handle the opposite case, $\frac{L_{b}\cdot\text{sup}_{\pi\in\phi_{b}}V^{\pi}(1_{h,x},P^{\prime})}{24HXA}>2\iota$. If the expected number of visits is high, we can use concentration inequalities to show that the probability of transitioning to an infrequent tuple must be very small. Denote $N_{h}^{\text{cru},b}(\mathcal{W}|(h,x,a))$ as the true times of entering $\mathcal{W}$ conditioned at time step $h$, visiting $(x_{h},a_{h})=(x,a)$. $\tilde{N}_{h}^{\text{cru},b}(\mathcal{W}|(h,x,a))$ denotes the private version. By Lemma 12 and a union bound, with high probability, for any $(x,a)\in\mathcal{X}\times\mathcal{A}$,
$6E_{\varepsilon,\delta}H^{2}X\iota$ $\geq\tilde{N}_{h}^{\text{cru},b}(\mathcal{W}|(h,x,a))\geq N_{h}^{\text{cru},b}(\mathcal{W}|(h,x,a))$ (15)
$\geq\frac{N^{\text{cru},b}_{h}(x,a)\mathbb{P}[\mathcal{W}|(h,x,a)]}{2}-\iota.$
The first inequality is due to the definition of $\mathcal{W}$, and the second inequality is because of our optimistic property of the private counter. Thus, this results in
$\mathbb{P}[\mathcal{W}|(h,x,a)]\leq\frac{672E_{\varepsilon,\delta}X^{2}AH^{3}\iota}{L_{b}\sup_{\pi\in\phi_{b}}V^{\pi}(1_{h,x},P^{\prime})}.$ (16)
In this case, we have,
$V^{\pi}(1_{h,x},P^{\prime})\max_{a\in\mathcal{A}}\mathbb{P}[\mathcal{W}|(h,x,a)]$ (17)
$\leq V^{\pi}(1_{h,x},P^{\prime})\cdot\frac{672E_{\varepsilon,\delta}X^{2}AH^{3}\iota}{L_{b}\sup_{\pi\in\phi_{b}}V^{\pi}(1_{h,x},P^{\prime})}$
$\leq\frac{672E_{\varepsilon,\delta}X^{2}AH^{3}\iota}{L_{b}}.$
Combining both cases,
$\sup_{\pi\in\phi_{b}}\mathbb{P}[\mathcal{B}_{h}|P,\pi]=\sup_{\pi\in\phi_{b}}\sum_{x\in\mathcal{X}}V^{\pi}(1_{h,x},P^{\prime})\max_{a\in\mathcal{A}}\mathbb{P}[\mathcal{W}|(h,x,a)]$ (18)
$\leq\sup_{\pi\in\phi_{b}}\sum_{x\in\mathcal{X}}\max\left(\frac{48HXA\iota}{L_{b}},\frac{672E_{\varepsilon,\delta}X^{2}AH^{3}\iota}{L_{b}}\right)$
$=\frac{672E_{\varepsilon,\delta}X^{3}AH^{3}\iota}{L_{b}},$
Summing over $H$ layers, we have with high probability,
$\sup_{\pi\in\phi_{b}}\mathbb{P}_{\pi}[\mathcal{B}]\leq\mathcal{O}\left(\frac{E_{\varepsilon,\delta}X^{3}AH^{4}\iota}{L_{b}}\right).$ (19)
To this end, we get the upper bound of the Model Bias term. For any reward function $r^{\prime}$, the non-negative inequality is obvious due to the “absorbing” MDP definition. For the right-hand side, we have that for any policy $\pi\in\phi_{b}$,
$V^{\pi}(r^{\prime},P)$ $=\sum_{S\notin\mathcal{B}}r^{\prime}(S)\mathbb{P}[S|P,\pi]+\sum_{S\in\mathcal{B}}r^{\prime}(S)\mathbb{P}[S|P,\pi]$
$=\sum_{S\notin\mathcal{B}}r^{\prime}(S)\mathbb{P}[S|P^{\prime},\pi]+\sum_{S\in\mathcal{B}}r^{\prime}(S)\mathbb{P}[S|P,\pi]$
$\leq V^{\pi}(r^{\prime},P^{\prime})+\sum_{S\in\mathcal{B}}H\mathbb{P}[S|P,\pi]$
$\leq V^{\pi}(r^{\prime},P^{\prime})+H\mathbb{P}[\mathcal{B}|P,\pi]$
$\leq V^{\pi}(r^{\prime},P^{\prime})+\frac{672E_{\varepsilon,\delta}X^{3}AH^{5}\iota}{L_{b}},$
where the inequality follows Lemma 16.
Proof of Lemma 9
This section provides the proof for Lemma 9, which bounds the difference between the value functions under the absorbing MDP ($P^{\prime}$) and the refined private estimate ($\widetilde{P}^{\text{ref},b}$). The proof first establishes that the fine exploration phase gathers sufficient data. It then introduces two key supporting lemmas regarding the uniform coverage of our fine exploration policy and the resulting visitation counts. These are then used in the main proof to bound the variance term.
A key prerequisite for this proof is that the fine exploration phase gathers sufficient visiting times for tuples not in $\mathcal{W}$. We establish this by leveraging the auxiliary policy $\pi_{0}$. By its construction in the crude exploration phase and the multiplicative Chernoff bound, we can show that for any tuple not in $\mathcal{W}$, running the auxiliary policy $\pi_{0}$ for $L_{b}$ episodes ensures its visitation count is sufficiently large (i.e., more than $C_{1}E_{\varepsilon,\delta}H^{2}\iota$). This ensures that the favorable properties of our model estimates, such as those established for the crude estimate in Lemma 14 and Lemma 15, also hold for the refined estimate, $\widetilde{P}^{\text{ref},b}$.
We then establish two supporting lemmas: a uniform coverage guarantee for our fine exploration policy $\pi^{\text{ref},b}$ and a high-probability bound on the resulting visitation counts.
Lemma 17
Conditioned on the event in Lemma 14, the policy $\pi^{\text{ref},b}$ in Algorithm 3 satisfies that $\max_{\mu\in\phi}\sum_{h,x,a}\frac{V^{\mu}(1_{h,x,a},P^{\prime})}{V^{\pi^{\text{ref},b}}(1_{h,x,a},P^{\prime})}\leq 12XAH.$
Proof:
For some policy $\mu$, it holds that $L.H.S\!\leq\!\sum_{h,x,a}\!\!\frac{3V^{\mu}(1_{h,x,a},\widetilde{P}^{\text{cru},b})}{\frac{1}{4}V^{\pi^{\text{ref},b}}(1_{h\!,x\!,a},\widetilde{P}^{\text{cru},b})}\!\leq\!\!12XAH,$ where the first inequality is due to the property of $\widetilde{P}^{\text{cru},b}$ in Lemma 15, and the last inequality is by plugging $\mathcal{X}=\{\{V^{\mu}(1_{h,\cdot,\cdot},\widetilde{P}^{\text{cru},b})\}_{h=1}^{H}\mid\mu\in\phi\},d=XA$ and $m=H$ into Lemma 13. ∎
Second, we provide a high-probability lower bound on the visitation counts collected by the fine exploration policy for all relevant state-action pairs.
Lemma 18
[…]
Proof:
The expectation visitation counts of $(h,x,a)$ from running $\pi^{\text{ref},b}$ for $L_{b}$ episodes is $L_{b}V^{\pi^{\text{ref},b}}(1_{h,x,a},P)\geq L_{b}V^{\pi^{\text{ref},b}}(1_{h,x,a},P^{\prime})$ by definition. According to Lemma 12, with probability $1-\frac{\delta}{HXAK}$, $N_{h}^{\text{ref},b}(x,a)\geq$ $\frac{1}{2}L_{b}V^{\pi^{\text{ref},b}}(1_{h,x,a},P)-\iota\geq\frac{1}{2}L_{b}V^{\pi^{\text{ref},b}}(1_{h,x,a},P^{\prime})-\iota$. The proof is done due to a union bond. ∎
Then we are ready to prove Lemma 9. For simplification, we define $\{f_{h}(\cdot)\}_{h=1}^{H}$ to be the value function of one policy $\mu$ under $P^{\prime}$ and $r^{\prime}$. Then it holds that, for any policy $\mu\in\phi_{b}$,
$|V^{\mu}\!(r^{\prime}\!,\!P^{\prime})-V^{\mu}\!(r^{\prime}\!,\!\widetilde{P}^{\text{ref},b})|$
$\leq$ $\sum_{h,x,a}\left|(\widetilde{P}^{\text{ref},b}_{h}-P^{\prime}_{h})f_{h+1}(x,a)\right|V^{\mu}(1_{h,x,a},\widetilde{P}^{\text{ref},b})$
$\leq$ $\!\sum_{h,x,a}\!\!\Bigg{(}\frac{N_{h}^{\text{ref},b}(x,a)}{\widetilde{N}_{h}^{\text{ref},b}(x,a)}\bigg{(}\!\sqrt{\frac{2\text{Var}_{P^{\prime}(\cdot|x,a)}f_{h+1}}{N_{h}^{\text{ref},b}(x,a)}}\!+\!\frac{2H\iota}{3N_{h}^{\text{ref},b}(x,a)}\!\bigg{)}$
$+\frac{2XHE_{\varepsilon,\delta}}{\widetilde{N}_{h}^{\text{ref},b}(x,a)}\Bigg{)}\cdot 4V^{\mu}(1_{h,x,a},P^{\prime})$
$\leq$ $10XHE_{\varepsilon,\delta}\iota\underbrace{\sum_{h,x,a}\frac{V^{\mu}(1_{h,x,a},P^{\prime})}{N_{h}^{\text{ref},b}(x,a)}}_{(i)}+4\sqrt{2\underbrace{\sum_{h,x,a}\frac{V^{\mu}(1_{h,x,a},P^{\prime})}{N_{h}^{\text{ref},b}(x,a)}}_{(i)}}$
$\cdot\sqrt{\underbrace{\sum_{h,x,a}V^{\mu}(1_{h,x,a},P^{\prime})\text{Var}_{P^{\prime}(\cdot|x,a)}f_{h+1}}_{(ii)}}$
$\leq$ $\mathcal{O}\bigg{(}\frac{X^{2}AH^{2}E_{\varepsilon,\delta}\iota}{L_{b}}+\sqrt{\frac{XAH^{3}\iota}{L_{b}}}\bigg{)}.$
The first inequality is because of the simulation lemma. The second inequality holds by the concentration of $\widetilde{P}^{\text{ref},b}$ by Bernstein inequality and Assumption 4, and Lemma 15. The third inequality is derived via Cauchy-Schwarz inequality, and the property of private counts. The last inequality comes from below.
The upper bound for $(i)$ is as follows.
$(i)$ $\leq\sum_{h,x,a}\frac{V^{\mu}(1_{h,x,a},P^{\prime})}{\max\{1,\frac{1}{2}L_{b}V^{\pi^{\text{ref},b}}(1_{h,x,a},P^{\prime})-\iota\}}$
$\leq$ $\sum_{h,x,a}V^{\mu}(1_{h,x,a},P^{\prime})\big{(}\mathbb{I}\{L_{b}V^{\pi^{\text{ref},b}}(1_{h,x,a},P^{\prime})<4\iota\}$
$+$ $\!\min\!\big{\{}1,\!\frac{4}{L_{b}\!V^{\pi^{\text{ref},b}}\!(1_{h,x,a},\!P^{\prime})}\big{\}}\mathbb{I}\{L_{b}V^{\pi^{\text{ref},b}}\!(1_{h,x,a},\!P^{\prime})\!\geq\!4\iota\}\big{)}$
$\leq$ $\sum_{h,x,a}\!\frac{V^{\mu}(1_{h,x,a},P^{\prime})}{V^{\pi^{\text{ref},b}}\!(1_{h,x,a},\!P^{\prime})}\!\cdot\!\frac{4\iota}{L_{b}}+\sum_{h,x,a}\!\frac{V^{\mu}(1_{h,x,a},P^{\prime})}{V^{\pi^{\text{ref},b}}\!(1_{h,x,a},P^{\prime})}\!\cdot\!\frac{4}{L_{b}}$
$\leq$ $\frac{48HXA}{L_{b}}+\frac{48HXA\iota}{L_{b}}.$
The first line is due to Lemma 18, and the second inequality uses the standard decomposition; the last follows Lemma 17.
The upper bound for $(ii)$ is shown below: $(ii)=\sum_{h,x,a}V^{\mu}(1_{h,x,a},P^{\prime})\text{Var}_{P^{\prime}(\cdot|x,a)}f_{h+1}(\cdot)\leq H^{2}$, where the inequality results from a recursive application of Law of Total Variance.
Proof of Lemma 10
Consider a batch dataset $D$ of $n$ users, we construct private counts $\widetilde{N}_{h}(x,a,x^{\prime}),\widetilde{N}_{h}(x,a),\widetilde{R}_{h}(x,a)$ for all $(h,x,a,x^{\prime})$ by using the proposed Privatizer in in Section V. We prove the two claims of Lemma 10 separately: the privacy guarantee and the utility guarantee.
Privacy guarantee. The proof follows three steps: the adapted shuffle binary summation mechanism is $(\varepsilon^{\prime},\beta^{\prime})$-SDP for one counter, then the Privatizer is $(\varepsilon,\beta)$-SDP for one batch users, and then the entire mechanism for all users is $B$-batch $(\varepsilon,\beta)$-SDP.
Firstly, we consider the shuffle binary mechanism with parameter $(\varepsilon^{\prime},\beta^{\prime})$. Consider two neighboring input $D=(0,d_{1},\dots,d_{n})$ and $D^{\prime}=(1,d_{1},\dots,d_{n})$, we define the random variable $Q$ to be the sum of all the random bits (i.e., $y$ or $y_{1},\dots,y_{m}$) over all users in $D$, and similarly define $Q^{\prime}$ with respect to $D^{\prime}$. We first claim that the output of the shuffler $(\mathcal{F}\circ\mathcal{E}^{n})(D)=Q+\sum_{i=1}^{n}d_{i}$ is $(\varepsilon^{\prime},\beta^{\prime})$-DP. We denote this random mechanism as $M^{*}(\cdot)$.
Since $Q$ is binomial in both regimes, by Chernoff bounds, for any $\beta^{\prime}>0$, it holds that $\mathbb{P}(\left|Q-\mathbb{E}[Q]\right|>\sqrt{3\mathbb{E}[Q]\log(2/\beta^{\prime})})\leq\beta^{\prime}$. Therefore, define confidence interval as $I_{c}=(\mathbb{E}[Q]-\sqrt{3\mathbb{E}[Q]\log(2/\beta^{\prime})},\mathbb{E}[Q]+\sqrt{3\mathbb{E}[Q]\log(2/\beta^{\prime})})$, and we get $\mathbb{P}(Q\notin I_{c})\leq\beta^{\prime}$. A similar result also applies for $Q^{\prime}$.
When $n\leq\tau$, $Q\sim\text{Binomial}(\lceil\frac{\tau}{n}\rceil\cdot n,1/2)$, and $\mathbb{E}[Q]=\lceil\frac{\tau}{n}\rceil\cdot\frac{n}{2}$. For any $q\in I_{c}$, and $\varepsilon^{\prime}<1$ it holds
$\frac{\mathbb{P}(Q=q)}{\mathbb{P}(Q^{\prime}=q-1)}=\frac{2\mathbb{E}[Q]-q+1}{q}$
$\leq\frac{\mathbb{E}[Q]+\sqrt{3\mathbb{E}[Q]\log(2/\beta^{\prime})}+1}{\mathbb{E}[Q]-\sqrt{3\mathbb{E}[Q]\log(2/\beta^{\prime})}}$
$\leq\frac{\tau/2+\sqrt{\tau/2\cdot 3\log(2/\beta^{\prime})}+1}{\tau/2-\sqrt{\tau/2\cdot 3\log(2/\beta^{\prime})}}$
$=\frac{1+\sqrt{6\log(2/\beta^{\prime})/\tau}+2/\tau}{1-\sqrt{6\log(2/\beta^{\prime})/\tau}}$
$=\frac{1+\varepsilon^{\prime}/4+2/\tau}{1-\varepsilon^{\prime}/4}\leq\frac{1+\varepsilon^{\prime}/4+\varepsilon^{\prime}/4}{1-\varepsilon^{\prime}/4}=\frac{1+\varepsilon^{\prime}/2}{1-\varepsilon^{\prime}/4}\leq e^{\varepsilon^{\prime}}.$
When $n>\tau$, $Q\sim\text{Binomial}(n,\frac{\tau}{2n})$, and $\mathbb{E}[Q]=\frac{\tau}{2}$. For any $q\in I_{c}$, and $\varepsilon^{\prime}<1$, it holds
$\frac{P(Q=q)}{P\left(Q^{\prime}=q-1\right)}=\frac{n-q+1}{q}\cdot\frac{\frac{\tau}{2n}}{1-\frac{\tau}{2n}}$
$\leq\frac{n-\tau/2+\sqrt{\frac{3}{2}\tau\log(2/\beta^{\prime})}+1}{\tau/2-\sqrt{\frac{3}{2}\tau\log(2/\beta^{\prime})}}\cdot\frac{\frac{\tau}{2n}}{1-\frac{\tau}{2n}}$
$=\frac{n-\tau/2+\sqrt{\frac{3}{2}\tau\log(2/\beta^{\prime})}+1}{n-\tau/2}\cdot\frac{\tau/2}{\tau/2-\sqrt{\frac{3}{2}\tau\log(2/\beta^{\prime})}}$
$=\left(1+\frac{\sqrt{\frac{3}{2}\tau\log(2/\beta^{\prime})}+1}{n-\tau/2}\right)\cdot\frac{1}{1-\sqrt{6\log(2/\beta^{\prime})/\tau}}$
$\leq\left(1+\sqrt{6\log(2/\beta^{\prime})/\tau}+2/\tau\right)\cdot\frac{1}{1-\sqrt{6\log(2/\beta^{\prime})/\tau}}$
$=\frac{1+\varepsilon^{\prime}/4+2/\tau}{1-\varepsilon^{\prime}/4}\leq\frac{1+\varepsilon^{\prime}/4+\varepsilon^{\prime}/4}{1-\varepsilon^{\prime}/4}\leq e^{\varepsilon^{\prime}}.$
Therefore, we can conclude that in both regimes of $n$, $\forall q\in I_{c},\frac{\mathbb{P}(Q=q)}{\mathbb{P}(Q^{\prime}=q-1)}\leq e^{\varepsilon^{\prime}}$, and similarly $\frac{\mathbb{P}(Q=q)}{\mathbb{P}(Q^{\prime}=q-1)}\geq e^{-\varepsilon^{\prime}}$.
Define $\Gamma=\sum_{j=1}^{n}d_{j}$ to be the true sum of $D$, and also be the true sum of $D^{\prime}$ minus one. Therefore, for any subset $E\subseteq\mathbb{N}$, we have
$\mathbb{P}[M^{*}(D)\in E]\leq\mathbb{P}[M^{*}(D)\in E\wedge Q\in I_{c}]+\mathbb{P}[Q\notin I_{c}]$ (20)
$\leq\sum_{s\in E}\mathbb{P}[M^{*}(D)=s\wedge Q\in I_{c}]+\beta$
$=\sum_{s\in E}\mathbb{P}[Q=s-\Gamma\wedge s-\Gamma\in I_{c}]+\beta^{\prime}$
$\leq e^{\varepsilon^{\prime}}\cdot\sum_{s\in E}\mathbb{P}[Q^{\prime}=s-\Gamma-1\wedge s-\Gamma\in I_{c}]+\beta^{\prime}$
$\leq e^{\varepsilon^{\prime}}\cdot\sum_{s\in E}\mathbb{P}[M^{*}(D^{\prime})=s]+\beta$
$=e^{\varepsilon^{\prime}}\cdot\mathbb{P}[M^{*}(D^{\prime})\in E]+\beta^{\prime}.$
A dual argument can also be obtained in a similar way, and then we can conclude $M^{*}$ (i.e., $(\mathcal{F}\circ\mathcal{E}^{n})$) is $(\varepsilon^{\prime},\beta^{\prime})$-DP.
Consider we are facing one batch users, for each $(h,x,a)$, we call the shuffle mechanism once. Using the standard composition theorem [5] over all $(h,x,a)$, and fact that one user difference only results in at most $H$ differences in these counters, similar to Lemma 34 of [33], the release of $\{\ddot{N}_{h}(x,a)\}_{h,x,a}$ satisfies $(\frac{\varepsilon}{3},\beta)$-SDP due to the choice of $(\varepsilon^{\prime},\beta^{\prime})$. Similarly, the release of $\{\ddot{N}_{h}(x,a)\}_{h,x,a,x^{\prime}}$ and $\{\widetilde{R}_{h}(x,a)\}_{h,x,a}$ also satisfy $(\frac{\varepsilon}{3},\beta)$-SDP. Then, by post-processing theorem and standard composition theorem in [5] the release of all private counts $\{\widetilde{N}_{h}(x,a)\}_{h,x,a},\{\widetilde{N}_{h}(x,a,x^{\prime})\}_{h,x,a,x^{\prime}},\{\widetilde{R}_{h}(x,a)\}_{h,x,a}$ satisfy $(\varepsilon,\beta)$-SDP. Therefore, by post-processing theorem, for all $B$ batches users, our Privatizer satisfies that $B$-batch $(\varepsilon,\beta)$-SDP.
[…]
Applying the confidence bound on each counter, we can establish the utility bound that with probability $1-3\delta$ for all $(h,x,a,x^{\prime})$,
$\left|\ddot{N}_{h}(x,a)-N_{h}(x,a)\right|\leq\tilde{\mathcal{O}}(\frac{H}{\varepsilon}),$ (21)
$\left|\ddot{N}_{h}(x,a,x^{\prime})-N_{h}(x,a,x^{\prime})\right|\leq\tilde{\mathcal{O}}(\frac{H}{\varepsilon}),$
$\left|\widetilde{R}_{h}(x,a)-R_{h}(x,a)\right|\leq\tilde{\mathcal{O}}(\frac{H}{\varepsilon}).$
Referring to the post-processing procedures in Section V, the shuffle Privatizer satisfies Assumption 4 with $E_{\varepsilon,\delta}=\tilde{\mathcal{O}}\left(\frac{H}{\varepsilon}\right)$, and $N_{h}(x,a)\leq\widetilde{N}_{h}(x,a)\leq N_{h}(x,a)+E_{\varepsilon,\delta}$. Furthermore, with the constraints of the optimization problem, we observe $\ddot{N}_{h}(x,a)=\sum_{x^{\prime}}\ddot{N}_{h}(x,a,x^{\prime})$, which also implies that $\widetilde{N}_{h}(x,a)=\sum_{x^{\prime}}\widetilde{N}_{h}(x,a,x^{\prime})$.